OpenStudy (anonymous):

convert 145 degrees into radians.

OpenStudy (anonymous):

\[\frac{31\pi }{ 36} or \frac{ 5\pi }{ 6 } or \frac{ 6\pi }{ 7 } or \frac{ 29\pi }{ 36}\]

OpenStudy (whpalmer4):

There are \(2\pi\) radians in 360 degrees. Can you figure out the proportion?

OpenStudy (whpalmer4):

That stuff you paste in from bagatrix is unreadable...

OpenStudy (anonymous):

what?

OpenStudy (whpalmer4):

@some_someone the post you deleted.

OpenStudy (anonymous):

i havent deleted any post?